# Discarding Lavrentiev’s Gap in Non-automonous and Non-Convex Variational Problems

### v1

@@ 2.3 Multidimensional examples
for which there is no Lavrentiev phenomenon according to Theorems 2, 3, and 4. For a quick summary, see Tables 1 and 2. We denote the integrand in (2.4) as $f=f(x,t,\xi)$, where $x$ stands for the spacial variable, the variable $t$ corresponds to the $u$-dependence of $f$, and the variable $\xi$ to its $\nabla u$-dependence. Let us emphasize that all examples with explicit $u$-dependence are novel, with the only remark that for $f=f(t,\xi)$ they follow from [Bousquet-Pisa] and for $f=f(x,t,\xi)$ with $(t,\xi)\mapsto f(x,t,\xi)$ being convex they are embraced by [BMT]. In this section we will use the following notation: for any real-valued function $w$ and a constant $C\in{\mathbb{R}}$, we write $C\ll w$ whenever there exists a constant $c\in{\mathbb{R}}$ such that $C<c\leq w(\cdot)$.
Examples built upon power models. The typical examples of functionals exhibiting the Lavrentiev phenomenon are those of the variable exponent and double phase growth, i.e., when the integrand in (2.4) is given by $\widetilde{f}_{\text{v}}(x,\xi)\coloneqq|\xi|^{\widetilde{r}(x)}$ and $\widetilde{f}_{\text{d}}(x,\xi)\coloneqq|\xi|^{p}+\widetilde{a}(x)|\xi|^{q}$, respectively. To exclude the Lavrentiev phenomenon for the functional involving $\widetilde{f}_{\text{v}}$, one usually requires that $1\ll\widetilde{r}\in L^{\infty}(\Omega)$ is log-Hölder continuous ($\widetilde{r}\in\mathcal{P}^{\text{log}}$), cf. [zh]. For the functional defined by $\widetilde{f}_{\text{d}}$, the typical assumptions are $1<p<q$ and $0\leq\widetilde{a}\in C^{0,{\varkappa}}$ with $q\leq p+{\varkappa}\max(1,p/N)$, see [zh, eslemi, bgs-arma]. The latter was recently improved in [bcdfm], allowing for $p$ and $q$ arbitrary far from each other, by considering in the place of $C^{0,{\varkappa}}$ a broader class of weights ${\mathcal{Z}}^{{\varkappa}}$, where ${\varkappa}\in(0,\infty)$ dictates a polynomial rate of vanishing decay of $\widetilde{a}$. We stress that upon these choices of parameters, the functions $\widetilde{f}_{\text{v}}$ and $\widetilde{f}_{\text{d}}$ satisfy condition ($H^{\rm iso}_{0}$) and the corresponding energy functionals are thus covered by Theorem 1.
@@
Due to the function $t\mapsto C(t)$ and the lack of concavity assumption on $\omega(\cdot,t)$, the condition (2.6) does not describe the modulus of continuity of $a$ with respect to $x$. In the special case of $\omega(s,t)=s^{\varkappa}$, the sub-class of ${\mathcal{Z}}_{\omega(\cdot,\cdot)}$ is ${\mathcal{Z}}^{{\varkappa}}$, which is meaningful for all ${\varkappa}\in(0,\infty)$. For more information about the class ${\mathcal{Z}}^{{\varkappa}}$ in the case when $a$ does not depend on $t$ we refer to [bcdfm].
@@
We observe that already mentioned examples embrace the best known conditions without $u$-dependence of the functional, cf. [eslemi, BCM, DeF-multi, badi, bcdfm, zh, bgs-arma]. At the same time, introducing this extra dependence does not complicate the conditions in an artificial way.
@@
$b(t)\left[\psi_{0}(t,|\xi|)+a(x,t)\psi_{1}(t,|\xi|)\right]$ $\forall\,t\ \psi_{0}(t,\cdot),\psi_{1}(t,\cdot)$$N$-function*, $a\in{\mathcal{Z}}_{\omega(\cdot,\cdot)}$
@@
Table 1: Main examples of integrands for functionals $\int_{\Omega}f(x,u,\nabla u){\,{\rm d}x}$ for which ($H^{\textrm{iso}}$) is satisfied and so for which we provide the absence of any Lavrentiev phenomenon by Theorem 4. We write $f=f(x,t,\xi)$, where $x$ stands for the spacial variable, variable $t$ corresponds to $u$-dependence of $f$, and variable $\xi$ to its $\nabla u$-dependence. In all examples in the table, $b\in C({\mathbb{R}},(0,+\infty))$, which might be relaxed due to Remark 2.3. The class ${\mathcal{Z}}_{\omega(\cdot,\cdot)}$ from (2.6) embraces (but is not restricted to) Hölder continuity; ${\mathcal{Z}}^{\varkappa}$, ${\varkappa}>0$ is its special case. Conditions on $\omega$ in the last example simplify under growth restrictions. *Assumption that $\psi_{0},\psi_{1}$ are $N$-functions is given for the simplicity of the exposition.
@@
To construct a more general example, we pick two increasing, non-negative, convex functions $\psi_{0}$ and $\psi_{1}$ on ${[0,\infty)}$ such that the function $\psi_{1}/\psi_{0}$ is non-decreasing. Upon setting
@@
Inspired by [Lukas23, Example 2.7 (2)] let us consider
@@
where $a\in{\mathcal{Z}}_{\omega(\cdot,\cdot)}\cap L^{\infty}(\Omega\times{\mathbb{R}})$ for $\omega(s)=\exp(-s^{-{\varkappa}})$. In [Lukas23, Example 2.7 (2)] the authors allow the parameter $q$ to be strictly less than $1$ so that $f_{\text{D}}$ is not convex in that case. We demand – via (2.7) – that ${\varkappa}>q\min(1,\tfrac{N}{p})$, which means that $p\geq 1$ and $q>0$ might be arbitrary if the function $a$ decays fast enough near the points where it vanishes. Moreover, unlike [Lukas23], for the absence of the Lavrentiev gap obtained via Theorem 3, we allow for functionals with explicit $u$-dependence (i.e. $a=a(x,t)$), see Example B.3 for details.
@@
[…]
@@
[…]
Table 2: Main examples of anisotropic integrands for functionals $\int_{\Omega}f(x,u,\nabla u){\,{\rm d}x}$ for which we provide the absence of the Lavrentiev phenomenon. We write $f=f(x,t,\xi)$, where $x$ stands for the spacial variable, variable $t$ corresponds to $u$-dependence of $f$, and variable $\xi$ to its $\nabla u$-dependence. In all examples in the table, $b\in C({\mathbb{R}},(0,+\infty))$, which might be relaxed due to Remark 2.3. The class ${\mathcal{Z}}_{\omega(\cdot,\cdot)}$ from (2.6) embraces (but is not restricted to) Hölder continuity; ${\mathcal{Z}}^{\varkappa}$, ${\varkappa}>0$ is its special case. *Assumption that certain functions are $N$-function is given for the simplicity of the exposition.
Fully anisotropic examples. Inspired by [BaaBy, DeFilippisRegularity, BaaByOh, DeF-multi] we present an anisotropic Orlicz multi-phase example. Let us consider radially increasing, convex functions $(\psi_{j})_{j=0}^{k}$, such that each $\psi_{j}$ with $1\leq j\leq k$ grows essentially faster than $\psi_{0}$ at infinity:
@@
As already observed in item (iii) before Theorem 3, to justify that this function satisfies ($H^{\textrm{conv}}$), it is enough that this fact holds for each ${\vartheta}_{j}$. Assuming further that the functions $\psi_{j}$ are isotropic, and relying on Theorem 5 below, we only need to require that condition ($H^{\textrm{iso}}$) is satisfied for every ${\vartheta}_{j}$ which is the case when $a_{j}\in{\mathcal{Z}}_{\omega_{j}(\cdot,\cdot)}$ for $\omega_{j}$ satisfying (2.7) with $\psi_{j}$ in the place of $\psi_{1}$. This condition applied to isotropic Lagrangian fully covers the scope of [BaaBy, Theorem 3.1] and extends it in three directions. We allow for explicit $u$-dependence of the considered functional and do not need to assume that $\psi_{j}\in\Delta_{2}$ for any $j$. Moreover, in [BaaBy] the imposed compatibility condition forces the closeness of phases expressed as $\limsup_{|\xi|\to\infty}\frac{\psi_{j}(x,|\xi|)}{|\xi|\psi_{0}(x,|\xi|)}<\infty$ for every $j$. Upon the condition above, the function $\omega_{j}$ does not play a role of the modulus of continuity, because we do not require its concavity, so no closeness of phases is needed.
@@ 3.1 Notation
$|\nabla u|>0\textrm{ a.e. on }\Lambda_{+}\textrm{ and }|\nabla u|=0\textrm{ a.e. on }\Lambda_{0}\,.$ (3.1)
@@
For a function $w:{\mathbb{R}}^{d}\to{\mathbb{R}}$ bounded from below by an affine function, we define its greatest convex minorant $w^{**}:{\mathbb{R}}^{d}\to{\mathbb{R}}$ as the supremum of all convex functions which are not larger than $w$ in the whole ${\mathbb{R}}^{d}$. It can also be obtained by applying the Young conjugation operation * twice, see e.g. [C-b, Corollary 2.1.42]. Note however that this latter fact will not be used in the sequel. We say that a function $w:{\mathbb{R}^{N}}\to{\mathbb{R}}$ is superlinear if
@@
We say that a function $w:[0,\infty)\to[0,\infty)$ is an $N$–function if it is convex, continuous, and such that $w(0)=0$, $\lim_{t\to 0}{w(t)}/{t}=0$ and $\lim_{t\to\infty}{w(t)}/{t}=\infty$.
@@
Another important tool is the disintegration of finite measures defined on $\Omega\times{\mathbb{R}}$. To detail this technique, we closely follow [AmFuPa, Section 2.5]. Let $\mu$ be a positive Radon measure on ${\mathbb{R}}$ and $(\nu_{t})_{t\in{\mathbb{R}}}$ a family of ${\mathbb{R}}^{N+1}$-valued measures on $\Omega$ such that the function $t\mapsto\nu_{t}(B)$ is $\mu$-measurable for any Borel set $B\subset\Omega$. We further assume that
@@
for every bounded Borel map $g:\Omega\times{\mathbb{R}}\to{[0,\infty)}$. The above definitions also make sense when $\nu_{t}$ is a positive finite measure. Then (3.9) holds true for any non-negative Borel function $g$. Conversely, given a finite ${\mathbb{R}}^{N+1}$-valued measure on $\Omega\times{\mathbb{R}}$, it can be written as a product of the form $\nu_{t}\otimes\mu$ as above, where $\mu$ is finite and $|\nu_{t}|(\Omega)=1$ for $\mu$ a.e. $t\in{\mathbb{R}}$, see [AmFuPa, Theorem 2.28].
@@ 3.2 Precise representatives
In this paragraph, we closely follow [MSZ]. Given $q\in{\mathbb{N}}$, a set $E\subset{\mathbb{R}}^{N}$ is said to be countably ${\mathscr{H}}^{q}$-rectifiable if there exist a family $(E_{k})_{k\in{\mathbb{N}}}$ of subsets in ${\mathbb{R}}^{q}$ and for every $k\in{\mathbb{N}}$, a Lipschitz map $f_{k}:E_{k}\to{\mathbb{R}}^{N}$ such that ${\mathscr{H}}^{q}\big{(}E\backslash\bigcup\limits_{k\in{\mathbb{N}}}f_{k}(E_{k})\big{)}=0$.
@@
[MSZ, Theorems 1.1 and 1.2] Let $u\in W^{1,1}(\Omega)$ be precisely represented. Then, the level set $u^{-1}(t)$ is countably ${\mathscr{H}}^{N-1}$-rectifiable for a.e. $t\in{\mathbb{R}}$, the graph ${\mathrm{Graph}_{u}}=\{(x,u(x)):x\in\Omega\}$ of $u$ is countably ${\mathscr{H}}^{N}$-rectifiable and moreover, for every measurable set $E\subset\Omega$, it holds
@@
for every $\phi\in C^{0}_{c}({\mathbb{R}}^{N})$. For a proof of this result, holding for any countably rectifiable set, see e.g. [Simon, Theorem 11.6].
@@ 4 Between balance conditions ($H^{\textrm{iso}}$) and ($H^{\textrm{conv}}$)
The arguments presented in this section enable to reduce the proofs of Theorems 4 and Theorem 1 to Theorem 3. This is a consequence of Theorem 5 below which can be seen as a counterpart of [ha-aniso, Theorem 1.2] for functions having isotropic or orthotropic structure. In both cases, it is shown that a balance between $f$ and $f^{-}_{B}$ on sub-level sets of $f^{-}_{B}$ given by ($H^{\textrm{iso}}$) implies a similar balance between $f$ and $(f^{-}_{B})^{**}$ as in ($H^{\textrm{conv}}$). The main difference between Theorem 5 and [ha-aniso, Theorem 1.2] is the position of the constant (we have it outside $f_{B}^{-}$ and $(f_{B}^{-})^{**}$, respectively). Our proof is based on essentially different geometrical observations.
Theorem 5.
  1. 1.
  2. 2.
    If $f$ admits an orthotropic decomposition in the sense of (2.3), with each $f_{i}$ satisfying ($H^{\textrm{iso}}$), then $f$ satisfies ($H^{\textrm{conv}}$).
We point out that the result above also allows reformulating [BMT, Hypothesis (H)] or [Lukas23, Assumption 1 (a4)] in the isotropic or orthotropic setting, by getting rid of the **∗ ∗ operator.
In the proof of Theorem 5, we need some properties of convex minorants that we now present.
@@
Let ${w}:[0,\infty)\to[0,\infty)$ be a continuous and non-decreasing function. For every $t\geq 0$, there exists $a_{t}\in[0,t]$ such that ${w}^{**}$ is affine on $[a_{t},t]$ and ${w}(a_{t})={w}^{**}(a_{t})$.
@@
Let us then assume that $w^{**}(t)<w(t)$ and that $w^{**}$ is unbounded. We take $a_{t}\leq t$ to be the minimal number such that $w^{**}$ is affine on $[a_{t},t]$. It suffices to show that $w^{**}(a_{t})=w(a_{t})$. The result holds if $a_{t}=0$. Hence, let us assume that $a_{t}>0$.
@@
As $w^{**}\leq\widetilde{w}$, and $w^{**}$ is the greatest convex minorant of $w$, we have $w^{**}=\widetilde{w}$. Therefore, $w^{**}$ is affine on $[s,s_{2}]$. This however means that $w^{**}$ is affine on $[s,t]$, where $s<a_{t}$, which contradicts the definition of $a_{t}$. Therefore, we have that $w^{**}(a_{t})=w(a_{t})$. ∎
The following lemma gives a lower bound on the derivative of ${w}^{**}$ whenever ${w}$ is the essential infimum of convex functions.
@@
Let $B$ be a non-empty Borel subset of ${\mathbb{R}}^{N}$ and $\big{\{}{w}_{y}\big{\}}_{y\in B}$ be a family of non-negative and non-decreasing convex functions on ${[0,\infty)}$. Let us denote ${w}\coloneqq\operatorname*{ess\,inf}_{y\in B}{w}_{y}$. Then for all $s\geq 0$, it holds that
@@
As the last inequality is true for a.e. $y\in B$, we get $\widetilde{w}\leq{w}$. Therefore, $\widetilde{w}$ is a convex minorant of ${w}$. However, as $\mathcal{d}_{s}>D^{+}{w}^{**}(s)$, we have ${w}^{**}(\tau)<\widetilde{w}(\tau)$ for some $\tau>s$, which is a contradiction. ∎
@@
For a.e. $y$, since ${w}_{y}$ is non-negative and vanishes at 00, it achieves its global minimum in 00. As ${w}_{y}$ is convex, this implies that it is non-decreasing. Therefore, ${w}$ is also non-decreasing as the essential infimum of non-decreasing functions. Consequently, ${w}^{**}$ is non-decreasing as well. Let us also notice that for a.e. $y$ and any $s\geq 0$, by convexity of ${w}_{y}$, it holds
@@
which means that the family $\{{w}_{y}\}_{y\in B(x,{\varepsilon})\cap\Omega}$ is uniformly locally Lipschitz, as by Lemma A.2, the function $f$ is bounded on bounded sets. Therefore, ${w}$ is continuous.
Let us now take $s\in{[0,\infty)}$ such that ${w}^{**}(s)+s^{\max(p,N)}\leq\mathcal{k}_{2}{\varepsilon}^{-N}$. As ${w}$ is non-decreasing and continuous, by Lemma 4.1, there exists $a_{s}\in[0,s]$ such that ${w}^{**}$ is affine on $[a_{s},s]$ and $w(a_{s})={w}^{**}(a_{s})$. For every $y\in B(x,{\varepsilon})$, we have $B(x,{\varepsilon})\subset B(y,2{\varepsilon})$, and thus
@@
where the last inequality follows from (4.1). In particular, the above inequality holds when $y=x$. Assume that $a_{s}<s$ and let $s^{\prime}\in(a_{s},s)$. Using that ${w}^{**}$ is affine on $[a_{s},s]$ together with  (4.2), we get
@@
Letting $\delta$ to 00 and then $s^{\prime}$ to $s$, we get ${w}^{**}(s)\geq\tfrac{1}{{\mathcal{C}}_{\widehat{\mathcal{k}}}}{w}(s)-1$, or equivalently,

### v3

@@ 2.2 Multidimensional examples
for which there is no Lavrentiev phenomenon according to Theorems 2, 3, and 5. For a quick summary, see Tables 1 and 2. We denote the integrand in (2.1) as $f=f(x,t,\xi)$, where $x$ stands for the spacial variable, the variable $t$ corresponds to the $u$-dependence of $f$, and the variable $\xi$ to its $\nabla u$-dependence. Let us emphasize that all examples with explicit $u$-dependence are novel, with the only remark that for $f=f(t,\xi)$ they follow from [16] and for $f=f(x,t,\xi)$ with $(t,\xi)\mapsto f(x,t,\xi)$ being convex they are embraced by [17]. In this section we will use the following notation: for any real-valued function $w$ and a constant $C\in{\mathbb{R}}$, we write $C\ll w$ whenever there exists a constant $c\in{\mathbb{R}}$ such that $C<c\leq w(\cdot)$.
Examples built upon power models. The typical examples of functionals exhibiting the Lavrentiev phenomenon are those of the variable exponent and double phase growth, i.e., when the integrand in (2.1) is given by $\widetilde{f}_{\text{v}}(x,\xi)\coloneqq|\xi|^{\widetilde{r}(x)}$ and $\widetilde{f}_{\text{d}}(x,\xi)\coloneqq|\xi|^{p}+\widetilde{a}(x)|\xi|^{q}$, respectively. To exclude the Lavrentiev phenomenon for the functional involving $\widetilde{f}_{\text{v}}$, one usually requires that $1\ll\widetilde{r}\in L^{\infty}(\Omega)$ is log-Hölder continuous ($\widetilde{r}\in\mathcal{P}^{\text{log}}$), cf. [49]. For the functional defined by $\widetilde{f}_{\text{d}}$, the typical assumptions are $1<p<q$ and $0\leq\widetilde{a}\in C^{0,{\varkappa}}$ with $q\leq p+{\varkappa}\max(1,p/N)$, see [49, 32, 18]. The latter was recently improved in [13], allowing for $p$ and $q$ arbitrary far from each other, by considering in the place of $C^{0,{\varkappa}}$ a broader class of weights ${\mathcal{Z}}^{{\varkappa}}$, where ${\varkappa}\in(0,\infty)$ dictates a polynomial rate of vanishing decay of $\widetilde{a}$. We stress that upon these choices of parameters, the functions $\widetilde{f}_{\text{v}}$ and $\widetilde{f}_{\text{d}}$ satisfy condition ($H^{\rm iso}_{0}$) and the corresponding energy functionals are thus covered by Theorem 1.
@@
Due to the function $t\mapsto C(t)$ and the lack of concavity assumption on $\omega(\cdot,t)$, the condition (2.3) does not describe the modulus of continuity of $a$ with respect to $x$. In the special case of $\omega(s,t)=s^{\varkappa}$, the sub-class of ${\mathcal{Z}}_{\omega(\cdot,\cdot)}$ is ${\mathcal{Z}}^{{\varkappa}}$, which is meaningful for all ${\varkappa}\in(0,\infty)$. For more information about the class ${\mathcal{Z}}^{{\varkappa}}$ in the case when $a$ does not depend on $t$ we refer to [13].
@@
We observe that already mentioned examples embrace the best known conditions without $u$-dependence of the functional, cf. [32, 14, 29, 7, 13, 49, 18]. At the same time, introducing this extra dependence does not complicate the conditions in an artificial way.
@@
$b(t)\left[\psi_{0}(t,|\xi|)+a(x,t)\psi_{1}(t,|\xi|)\right]$ $\forall\,t\ \psi_{0}(t,\cdot),\psi_{1}(t,\cdot)$NN-function*, $a\in{\mathcal{Z}}_{\omega(\cdot,\cdot)}$
@@
Table 1: Main examples of integrands for functionals $\int_{\Omega}f(x,u,\nabla u){\,{\rm d}x}$ for which ($H^{\textrm{iso}}$) is satisfied and so for which we provide the absence of any Lavrentiev phenomenon by Theorem 5. We write $f=f(x,t,\xi)$, where $x$ stands for the spacial variable, variable $t$ corresponds to $u$-dependence of $f$, and variable $\xi$ to its $\nabla u$-dependence. In all examples in the table, $b\in C({\mathbb{R}},(0,+\infty))$, which might be relaxed due to Remark 2.1. The class ${\mathcal{Z}}_{\omega(\cdot,\cdot)}$ from (2.3) embraces (but is not restricted to) Hölder continuity; ${\mathcal{Z}}^{\varkappa}$, ${\varkappa}>0$ is its special case. Conditions on $\omega$ in the last example simplify under growth restrictions. *Assumption that $\psi_{0},\psi_{1}$ are NN-functions is given for the simplicity of the exposition.
@@
To construct a more general example, we pick two increasing, non-negative, convex functions $\psi_{0}$ and $\psi_{1}$ on $0,\infty)$} such that the function $\psi_{1}/\psi_{0}$ is non-decreasing. Upon setting
@@
Inspired by [43, Example 2.7 (2)] let us consider
@@
where $a\in{\mathcal{Z}}_{\omega(\cdot,\cdot)}\cap L^{\infty}(\Omega\times{\mathbb{R}})$ for $\omega(s)=\exp(-s^{-{\varkappa}})$. In [43, Example 2.7 (2)] the authors allow the parameter $q$ to be strictly less than $1$ so that $f_{\text{D}}$ is not convex in that case. We demand – via (2.4) – that ${\varkappa}>q\min(1,\tfrac{N}{p})$, which means that $p\geq 1$ and $q>0$ might be arbitrary if the function $a$ decays fast enough near the points where it vanishes. Moreover, unlike [43], for the absence of the Lavrentiev gap obtained via Theorem 3, we allow for functionals with explicit $u$-dependence (i.e. $a=a(x,t)$), see Example B.3 for details.
@@
$b(t)\left[\psi_{0}(|\xi|)+\sum_{i=1}^{N}a_{i}(x,t)\psi_{i}(|\xi_{i}|)\right]$ $\forall\,i\ \psi_{i}$NN-function*, $\psi_{i}/\psi_{0}$ – non-decreasing Theorem 3
@@
$b(t)\left(\psi(|\langle{\upsilon}(x),\xi\rangle|)+|\xi|^{N/\gamma}\right)$ $\psi\in\Delta_{2}$NN-function*, ${\upsilon}\in C^{0,\gamma}(\Omega;{\mathbb{R}^{N}})$, $\gamma\in(0,1]$ Theorem 2
Table 2: Main examples of anisotropic integrands for functionals $\int_{\Omega}f(x,u,\nabla u){\,{\rm d}x}$ for which we provide the absence of the Lavrentiev phenomenon. We write $f=f(x,t,\xi)$, where $x$ stands for the spacial variable, variable $t$ corresponds to $u$-dependence of $f$, and variable $\xi$ to its $\nabla u$-dependence. In all examples in the table, $b\in C({\mathbb{R}},(0,+\infty))$, which might be relaxed due to Remark 2.1. The class ${\mathcal{Z}}_{\omega(\cdot,\cdot)}$ from (2.3) embraces (but is not restricted to) Hölder continuity; ${\mathcal{Z}}^{\varkappa}$, ${\varkappa}>0$ is its special case. *Assumption that certain functions are NN-function is given for the simplicity of the exposition.
Fully anisotropic examples. Inspired by [4, 30, 5, 29] we present an anisotropic Orlicz multi-phase example. Let us consider radially increasing, convex functions $(\psi_{j})_{j=0}^{k}$, such that each $\psi_{j}$ with $1\leq j\leq k$ grows essentially faster than $\psi_{0}$ at infinity:
@@
As already observed in item (iii) before Theorem 3, to justify that this function satisfies ($H^{\textrm{conv}}$), it is enough that this fact holds for each ${\vartheta}_{j}$. Assuming further that the functions $\psi_{j}$ are isotropic, and relying on Theorem 4 below, we only need to require that condition ($H^{\textrm{iso}}$) is satisfied for every ${\vartheta}_{j}$ which is the case when $a_{j}\in{\mathcal{Z}}_{\omega_{j}(\cdot,\cdot)}$ for $\omega_{j}$ satisfying (2.4) with $\psi_{j}$ in the place of $\psi_{1}$. This condition applied to isotropic Lagrangian fully covers the scope of [4, Theorem 3.1] and extends it in three directions. We allow for explicit $u$-dependence of the considered functional and do not need to assume that $\psi_{j}\in\Delta_{2}$ for any $j$. Moreover, in [4] the imposed compatibility condition forces the closeness of phases expressed as $\limsup_{|\xi|\to\infty}\frac{\psi_{j}(x,|\xi|)}{|\xi|\psi_{0}(x,|\xi|)}<\infty$ for every $j$. Upon the condition above, the function $\omega_{j}$ does not play a role of the modulus of continuity, because we do not require its concavity, so no closeness of phases is needed.
@@ 3.1 Notation
[…]
@@
For a function $w:{\mathbb{R}}^{d}\to{\mathbb{R}}$ bounded from below by an affine function, we define its greatest convex minorant $w^{**}:{\mathbb{R}}^{d}\to{\mathbb{R}}$ as the supremum of all convex functions which are not larger than $w$ in the whole ${\mathbb{R}}^{d}$. It can also be obtained by applying the Young conjugation operation * twice, see e.g. [23, Corollary 2.1.42]. Note however that this latter fact will not be used in the sequel. We say that a function $w:{\mathbb{R}^{N}}\to{\mathbb{R}}$ is superlinear if
@@
We say that a function $w:[0,\infty)\to[0,\infty)$ is an NN–function if it is convex, continuous, and such that $w(0)=0$, $\lim_{t\to 0}{w(t)}/{t}=0$ and $\lim_{t\to\infty}{w(t)}/{t}=\infty$.
@@
Another important tool is the disintegration of finite measures defined on $\Omega\times{\mathbb{R}}$. To detail this technique, we closely follow [2, Section 2.5]. Let $\mu$ be a positive Radon measure on $\mathbb{R}$} and $(\nu_{t})_{t\in{\mathbb{R}}}$ a family of ${\mathbb{R}}^{N+1}$-valued measures on $\Omega$ such that the function $t\mapsto\nu_{t}(B)$ is $\mu$-measurable for any Borel set $B\subset\Omega$. We further assume that
@@
for every bounded Borel map $g:\Omega\times{\mathbb{R}}\to{[0,\infty)}$. The above definitions also make sense when $\nu_{t}$ is a positive finite measure. Then (3.9) holds true for any non-negative Borel function $g$. Conversely, given a finite ${\mathbb{R}}^{N+1}$-valued measure on $\Omega\times{\mathbb{R}}$, it can be written as a product of the form $\nu_{t}\otimes\mu$ as above, where $\mu$ is finite and $|\nu_{t}|(\Omega)=1$ for $\mu$ a.e. $t\in{\mathbb{R}}$, see [2, Theorem 2.28].
@@ 3.2 Precise representatives
[…]
@@
[45, Theorems 1.1 and 1.2] Let $u\in W^{1,1}(\Omega)$ be precisely represented. Then, the level set $u^{-1}(t)$ is countably ${\mathscr{H}}^{N-1}$-rectifiable for a.e. $t\in{\mathbb{R}}$, the graph ${\mathrm{Graph}_{u}}=\{(x,u(x)):x\in\Omega\}$ of $u$ is countably ${\mathscr{H}}^{N}$-rectifiable and moreover, for every measurable set $E\subset\Omega$, it holds
@@
for every $\phi\in C^{0}_{c}({\mathbb{R}}^{N})$. For a proof of this result, holding for any countably rectifiable set, see e.g. [48, Theorem 11.6].
@@ 4 Between balance conditions ($H^{\textrm{iso}}$) and ($H^{\textrm{conv}}$)
This section is devoted to the proof of Theorem 4. The reasoning is based on several geometrical observations of the properties of convex minorants, which we now present.
@@
Let ${w}:[0,\infty)\to[0,\infty)$ be a continuous and non-decreasing function. For every $t\geq 0$, there exists $a_{t}\in[0,t]$ such that ${w}^{**}$ is affine on $a_{t},t]$ and ${w}(a_{t})={w}^{**}(a_{t})$.
@@
Let us then assume that $w^{**}(t)<w(t)$ and that $w^{**}$ is unbounded. We take $a_{t}\leq t$ to be the minimal number such that $w^{**}$ is affine on $a_{t},t]$. It suffices to show that $w^{**}(a_{t})=w(a_{t})$. The result holds if $a_{t}=0$. Hence, let us assume that $a_{t}>0$.
@@
As $w^{**}\leq\widetilde{w}$, and $w^{**}$ is the greatest convex minorant of $w$, we have $w^{**}=\widetilde{w}$. Therefore, $w^{**}$ is affine on $s,s_{2}]$. This however means that $w^{**}$ is affine on $[s,t]$, where $s<a_{t}$, which contradicts the definition of $a_{t}$. Therefore, we have that $w^{**}(a_{t})=w(a_{t})$. ∎
The following lemma gives a lower bound on the derivative of ${w}^{**}$ whenever w{w} is the essential infimum of convex functions.
@@
Let $B$ be a non-empty Borel subset of ${\mathbb{R}}^{N}$ and $\big\{{w}_{y}\big\}_{y\in B}$ be a family of non-negative and non-decreasing convex functions on $0,\infty)$}. Let us denote ${w}\coloneqq\operatorname*{ess\,inf}_{y\in B}{w}_{y}$. Then for all $s\geq 0$, it holds that
@@
As the last inequality is true for a.e. $y\in B$, we get $\widetilde{w}\leq{w}$. Therefore, $\widetilde{w}$ is a convex minorant of w{w}. However, as $\mathcal{d}_{s}>D^{+}{w}^{**}(s)$, we have ${w}^{**}(\tau)<\widetilde{w}(\tau)$ for some $\tau>s$, which is a contradiction. ∎
@@
For a.e. $y$, since ${w}_{y}$ is non-negative and vanishes at 0, it achieves its global minimum in 0. As ${w}_{y}$ is convex, this implies that it is non-decreasing. Therefore, w{w} is also non-decreasing as the essential infimum of non-decreasing functions. Consequently, ${w}^{**}$ is non-decreasing as well. Let us also notice that for a.e. $y$ and any $s\geq 0$, by convexity of ${w}_{y}$, it holds
@@
which means that the family $\{{w}_{y}\}_{y\in B(x,{\varepsilon})\cap\Omega}$ is uniformly locally Lipschitz, as by Lemma A.2, the function $f$ is bounded on bounded sets. Therefore, w{w} is continuous.
Let us now take $s\in{[0,\infty)}$ such that ${w}^{**}(s)+s^{\max(p,N)}\leq\mathcal{k}_{2}{\varepsilon}^{-N}$. As w{w} is non-decreasing and continuous, by Lemma 4.1, there exists $a_{s}\in[0,s]$ such that ${w}^{**}$ is affine on $a_{s},s]$ and $w(a_{s})={w}^{**}(a_{s})$. For every $y\in B(x,{\varepsilon})$, we have $B(x,{\varepsilon})\subset B(y,2{\varepsilon})$, and thus
@@
where the last inequality follows from (4.1). In particular, the above inequality holds when $y=x$. Assume that $a_{s}<s$ and let $s^{\prime}\in(a_{s},s)$. Using that ${w}^{**}$ is affine on $a_{s},s]$ together with  (4.2), we get
@@
Letting $\delta$ to 0 and then $s^{\prime}$ to $s$, we get ${w}^{**}(s)\geq\tfrac{1}{{\mathcal{C}}_{\widehat{\mathcal{k}}}}{w}(s)-1$, or equivalently,